You have $10,000 to invest in a stock portfolio. Your choices are Stock X with an expected return of 21.0% and a standard deviation of 40% and T-Bills (e.g., the risk free asset) with an expected return of 5% and a standard deviation of 0%. How much money will you invest in Stock X if your goal is to create a portfolio with an expected return of 26%?
Answer #1

Portfolio Ret = Weighted Avg ret of securities in that portfolio.

Let X be the weight of stock A & 1-X is weight of stock B.

Particulars Weight Ret Wtd Ret
X X 0.21 0.21X
Y 1 - X 0.05 0.05 - 0.05X
Portfolio Ret 0.16X + 0.05

Thus 0.16X + 0.05 = 0.26

0.16X = 0.26-0.05

= 0.21

X = 0.21 / 0.16

= 1.3125

Amount invested in Stock X = 1.3125 * 10000

= $ 13125

Pls comment, if any further assistance is required